# Entanglement and Bell Nonlocality in $\tau^{+}\tau^{-}$ at the LHC using Machine Learning for Neutrino Reconstruction

### v1

@@ 1 Introduction
In addition to the $t\bar{t}$ system, there has been early work on a number of other final states, including $WW$ Aguilar-Saavedra:2022mpg ; Fabbrichesi:2023cev ; Morales:2023gow ; Bi:2023uop , $ZZ$ Aguilar-Saavedra:2022wam ; Fabbrichesi:2023cev ; Morales:2023gow ; Bernal:2023ruk ; Ruzi:2024cbt ; Wu:2024ovc , $\Lambda_{b}\bar{\Lambda}_{b}$ Afik:2024uif , $Y\bar{Y}$ Wu:2024asu , $\phi\phi$ Gabrielli:2024kbz , $\mu^{+}\mu^{-}$ Gao:2024leu , and $q\bar{q}$ Cheng:2025cuv . Some of these other states go beyond two qubits and probe qutrits, multipartite systems Sakurai:2023nsc ; Aguilar-Saavedra:2024whi ; Morales:2024jhj ; Subba:2024mnl , and other quantum informational concepts Altomonte:2023mug ; Barr:2024djo ; Altomonte:2024upf . Studies of the $\tau^{+}\tau^{-}$ system have been suggested at $e^{+}e^{-}$ colliders, including at the FCC-ee Altakach:2022ywa ; Fabbrichesi:2024wcd , at the CEPC Ma:2023yvd , at Belle-II Ehataht:2023zzt , and at the BEPC Han:2025ewp .
Generally, $t\bar{t}$ has been considered the canonical qubit system in hadron colliders, whereas $\tau^{+}\tau^{-}$ has been considered the canonical qubit system in lepton colliders. Although both systems have at least two neutrinos in their final states, in the $t\bar{t}$ system there are enough kinematic constraints, save for discrete ambiguities, to solve the neutrino momenta at hadron or lepton colliders, while for $\tau^{+}\tau^{-}$ there are only enough constraints at lepton colliders due to the well-determined collision energies. Because of this difficulty, there has not been a study of $\tau^{+}\tau^{-}$ state at the LHC. [^1] Ref. Fabbrichesi:2022ovb presented theoretical-level results for $\tau^{+}\tau^{-}$ at the LHC but did not provide a practical way to measure the quantum state. In this work, we leverage the power of ML to open up this notoriously difficult channel and show that it has excellent sensitivity to detecting quantum entanglement and Bell nonlocality, already with the LHC’s current dataset.
@@ 2.2 The Decay Method
The spin analyzing power varies from $-1$ to $+1$ corresponding to maximally anti-correlated and maximally correlated, respectively. A spin analyzing power of 00 indicates no correlation between the spin and the spin analyzer direction.
@@ 2.4 The $\tau^{+}\tau^{-}$ State
The quantum state of $\tau^{+}\tau^{-}$ in production $pp\to\tau^{+}\tau^{-}X$ has been discussed in Ref. Fabbrichesi:2022ovb . Below the $Z$ boson mass $m_{Z}$ where the $Z$ contribution is negligible, the $\tau^{+}\tau^{-}$ state is dominantely produced through an $s$-channel photon leading to a triplet state that is nearly maximally entangled. As the $Z$ contribution grows and the interference with the $\gamma$ contribution becomes sizable, the mixing between the $Z$ and $\gamma$ leads to a nearly separable state as seen in Fig. 1. The minimum entanglement occurs near $m_{\tau\tau}\approx 70~{}{\rm GeV}$.
@@ 4 Neutrino Momentum Reconstruction Using a Generative Model
Compared with existing approaches, our method presents several notable advancements. Firstly, we introduce an additional Transformer layer within the generation head, thereby significantly enhancing the representational capability of the model. Secondly, we optimize the sampling procedure by replacing the modified Denoising Diffusion Probabilistic Model (DDPM) used in OmniLearn with the simpler Denoising Diffusion Implicit Model (DDIM), enabling increased sampling steps without incurring substantial computational costs. Thirdly, we incorporate random rotations into the training data augmentation process, greatly enlarging the dataset and enabling larger batch sizes, thus improving the training efficiency of the Transformer model while explicitly leveraging the rotational symmetry inherent in physical processes. Lastly, the nearest-neighbor calculations are dynamically tailored according to the number of visible components resulting from $\tau$ decays; specifically, we select the six nearest neighbors for decay channels involving rho particles and only two nearest neighbors for other channels, effectively capturing the distinct physical characteristics associated with each decay channel.
@@ 4.1 Model Architecture and Training Strategy
A key limitation of generative approaches in collider physics is their reliance on large Monte Carlo datasets. To mitigate this, we employ an efficient data augmentation strategy, where each event is randomly rotated in the transverse plane relative to the center-of-mass frame. This technique enhances generalization without requiring additional MC event generation, significantly reducing computational costs. By applying this augmentation, we increase the training statistics by approximately 50 times, reaching a range of $1$M to $10$M effective events across all subchannels.
@@ 5.1 Event Selection and Categorization
This study focuses on events where $\tau$ decays result in a single charged track, as these modes have a higher branching ratio and greater spin analyzing power compared to multi-track decays. Since $\tau$ decays involve neutrinos and various hadronic final states, the spin correlations between the two $\tau$s are encoded in their decay products. To preserve this information, events are categorized into distinct subchannels based on their reconstructed final states, each exhibiting a different spin analyzing power.
At the reconstruction level, events are categorized into seven subchannels: three hadronic-hadronic subchannels, where both $\tau$s decay hadronically, and four leptonic-hadronic subchannels, where one $\tau$ decays leptonically while the other undergoes a hadronic decay. The classification of hadronic $\tau$ decays is based on prongness, which is inherently a reconstruction-level concept rather than a direct representation of the true decay products. In practice, the detailed composition of the $\tau$ decay cannot be fully reconstructed; instead, hadronic $\tau$s are identified based on the number of detected charged tracks and associated neutral energy deposits in the calorimeter. A one-prong (1p) $\tau$ corresponds to a single reconstructed charged track. The neutral component is inferred from calorimeter clusters associated with neutral pions. Consequently, the notation “XpYn” refers to the reconstructed prongness, where X denotes the number of charged tracks, and Y represents the number of neutral clusters attributed to the $\tau$ decay. The classification of subchannels, along with their corresponding spin-analyzing power, is summarized in Table 5.
To ensure a clean selection of $pp\to Z/\gamma^{*}\to\tau^{+}\tau^{-}$ events, a set of preselection criteria is applied. The analysis is performed using $\tau$ candidates, which represent the visible components of $\tau$ decays, while the invisible component, consisting of neutrinos, is reconstructed in a later stage. Events containing $b$-tagged jets are removed to suppress backgrounds from top-quark decays. For hadronic-hadronic subchannels, events with reconstructed electrons or muons are vetoed to avoid contamination from leptonic $\tau$ decays or misidentified leptons. Exactly two $\tau$ candidates are required, with opposite electric charges, ensuring a well-defined event topology. The classification of each event is then determined based on the reconstructed decay mode of the $\tau$s. Additional kinematic requirements are imposed, requiring transverse momentum $p_{T}>10$ GeV and pseudorapidity $|\eta|<2.5$ for the $\tau$ candidates. The full reconstruction of the $\tau$s, including the neutrino kinematics, is addressed in Sec. 4. For the final signal region selection, events must satisfy the invariant mass and center-of-mass $\tau$ angle constraints of
$1.080~{}{\rm GeV}<m_{\tau\tau}<100~{}{\rm GeV},\qquad\qquad 0.6<\theta_{\tau}/2\pi<$ 1.0 . (19)
@@
Subchannel Prongness $|\kappa_{A}\times\kappa_{B}|$ $\tau\tau\to\ell\pi$ $\tau\tau\to\pi\rho$ $\tau\tau\to\pi\pi$ $\tau\tau\to\ell\rho$ $\tau\tau\to\rho\rho$
@@
$e\pi$ e - 1p0n 0.33 $1245.85\pm 3.25$ $80.33\pm 1.23$ $0.09\pm 0.02$ $0.14\pm 0.03$ $0.01\pm 0.01$
$\mu\pi$ $\mu$ 1p0n 0.34 $1640.89\pm 3.73$ $106.80\pm 1.42$ $0.04\pm 0.01$ $0.05\pm 0.02$ $<0.01$
$e\rho$ e - 1p1n 0.14 $25.32\pm 0.46$ $3603.19\pm 8.23$ $<0.01$ $0.13\pm 0.03$ $0.47\pm 0.08$
$\mu\rho$ $\mu$ 1p1n 0.15 $36.45\pm 0.56$ $4869.69\pm 9.57$ $<0.01$ $0.09\pm 0.02$ $0.31\pm 0.06$
Signal Region ($80~{}{\rm GeV}<m_{\tau\tau}<100~{}{\rm GeV}$ $0.6<\theta_{\tau}/2\pi<1)$
@@ 5.2 Backgrounds and Systematic Uncertainties
Top-quark backgrounds arise primarily from $t\bar{t}$ production, where both top quarks decay via $t\to Wb$, and the subsequent $W\to\tau\nu$ decay can produce prompt $\tau$s. This background includes genuine $\tau^{+}\tau^{-}$ pairs but is largely suppressed by requiring events to be free of $b$-tagged jets.
@@
$\pi\pi$ $<0.01$ $0.41\pm 0.29$ $<0.01$ $0.96\pm 0.22$ $<0.01$ $1.37\pm 0.37$
$\pi\rho$ $0.77\pm 0.55$ $1.45\pm 0.55$ $<0.01$ $3.28\pm 0.41$ $<0.01$ $5.50\pm 0.87$
$\rho\rho$ $<0.01$ $4.14\pm 0.93$ $<0.01$ $10.13\pm 0.71$ $<0.01$ $14.27\pm 1.17$
$e\pi$ $4.25\pm 1.28$ $3.31\pm 0.83$ $1.51\pm 0.75$ $42.33\pm 1.46$ $<0.01$ $51.40\pm 2.24$
$e\rho$ $28.20\pm 3.30$ $12.43\pm 1.60$ $2.64\pm 1.00$ $120.53\pm 2.46$ $<0.01$ $163.80\pm 4.53$
$\mu\pi$ $2.70\pm 1.02$ $5.39\pm 1.06$ $<0.01$ $50.89\pm 1.60$ $<0.01$ $58.98\pm 2.17$
$\mu\rho$ $34.00\pm 3.62$ $12.84\pm 1.63$ $<0.01$ $<0.01$ $<0.01$ $46.84\pm 3.97$
Signal Region ($80~{}{\rm GeV}<m_{\tau\tau}<100~{}{\rm GeV}$ $0.6<\theta_{\tau}/2\pi<1)$
$\pi\pi$ $<0.01$ $<0.01$ $<0.01$ $0.05\pm 0.05$ $<0.01$ $0.05\pm 0.05$
$\pi\rho$ $<0.01$ $<0.01$ $<0.01$ $0.20\pm 0.10$ $<0.01$ $0.20\pm 0.10$
$\rho\rho$ $<0.01$ $<0.01$ $<0.01$ $0.96\pm 0.22$ $<0.01$ $0.96\pm 0.22$
$e\pi$ $<0.01$ $0.21\pm 0.21$ $<0.01$ $3.93\pm 0.45$ $<0.01$ $4.14\pm 0.49$
$e\rho$ $2.70\pm 1.02$ $0.62\pm 0.36$ $<0.01$ $12.60\pm 0.80$ $<0.01$ $15.92\pm 1.34$
$\mu\pi$ $<0.01$ $0.83\pm 0.41$ $<0.01$ $5.59\pm 0.53$ $<0.01$ $6.42\pm 0.67$
$\mu\rho$ $2.32\pm 0.95$ $1.04\pm 0.46$ $<0.01$ $<0.01$ $<0.01$ $3.35\pm 1.05$
@@ 5.2.1 Systematics Uncertainties
Object-level systematic uncertainties are implemented by modifying the kinematics of reconstructed objects before neutrino reconstruction, simulating realistic detector effects. Following the standard approach Cowan:2010js , uncertainties are varied by ±$\pm 1\sigma$. The jet energy scale (JES) and tau energy scale (TES) are adjusted by 5% ATLAS-JET-PERF and 3% ATLAS-TAU-PERF , respectively, directly impacting the transverse momenta of jets and $\tau$s, which in turn modify the missing transverse energy (MET). Additionally, two independent soft MET components, each with a 1 GeV variation in the $x$ and $y$ directions, are introduced to account for potential soft energy fluctuations in the detector.
@@
Figure 4: Comparison of reconstructed neutrino kinematic distributions under systematic variations from soft MET along the $x$ direction for $\ell\rho$ channel. The three columns correspond to transverse momentum ($p_{T}$), pseudorapidity ($\eta$), and azimuthal angle ($\phi$). Each plot includes the nominal reconstruction (teal) along with systematically shifted variations: $+1\sigma$ (goldenrod) and $-1\sigma$ (rose) shifts applied to neutrinos originating from $\tau^{+}$. The middle panels show the ratio of systematically varied reconstructions and nominal values to the truth values, while the bottom panels contain two-dimensional correlation contour plots illustrating the linear dependency between the truth and systematically shifted predictions. The contour lines represent density levels, with numerical labels indicating the percentage of total data enclosed within each contour.
@@ 5.3 Trigger Considerations
The di-tau trigger222 HLT_tau35_medium1_tracktwo_tau25_medium1_tracktwo_L1TAU20IM_2TAU12IM is designed for events with two hadronically decaying $\tau$-leptons, covering the $\pi\pi$, $\pi\rho$, and $\rho\rho$ subchannels. This trigger requires the leading $\tau$ to have $p_{T}>35$ GeV and the subleading $\tau$ to have $p_{T}>25$ GeV. The tau + muon trigger333 HLT_mu14_tau25_medium1_tracktwo is applied to events in the $\mu\pi$ and $\mu\rho$ channels, requiring a $\tau$ with $p_{T}>25$ GeV and a muon with $p_{T}>14$ GeV. Additionally, the single-muon trigger444 HLT_mu26_ivarmedium is used, requiring a muon with $p_{T}>26$ GeV. For events involving electrons, the tau + electron trigger555 HLT_e17_lhmedium_nod0_ivarloose_tau25_medium1_tracktwo is applied, requiring a $\tau$ with $p_{T}>25$ GeV and an electron with $p_{T}>17$ GeV. The single electron trigger666 HLT_e26_lhtight_nod0_ivarloose is also utilized, selecting events with an electron satisfying $p_{T}>26$ GeV.
To emulate the trigger effect in our study, we impose these trigger conditions in addition to the signal region selection criteria. The di-tau trigger imposes a significant efficiency penalty on events with low-$p_{T}$ visible $\tau$s, particularly for single-pion decays. However, both the $\rho\rho$ and $\pi\rho$ channels remain viable under the di-tau trigger. In contrast, the tau + muon trigger accommodates both the $\mu\pi$ and $\mu\rho$ channels, leading to a higher event rate. Similarly, the tau + electron trigger allows for the $e\pi$ and $e\rho$ channels, expanding the accessible phase space. Additionally, the single-lepton triggers for muons and electrons provide an independent selection pathway for events containing a high-$p_{T}$ isolated lepton.

### v2

@@ 1 Introduction
In addition to the $t\bar{t}$ system, there has been early work on a number of other final states, including $WW$ Aguilar-Saavedra:2022mpg ; Fabbrichesi:2023cev ; Morales:2023gow ; Bi:2023uop ; Grossi:2024jae , $ZZ$ Aguilar-Saavedra:2022wam ; Fabbrichesi:2023cev ; Morales:2023gow ; Bernal:2023ruk ; Ruzi:2024cbt ; Wu:2024ovc ; Grossi:2024jae , $\Lambda_{b}\bar{\Lambda}_{b}$ Afik:2024uif , $Y\bar{Y}$ Wu:2024asu , $\phi\phi$ Gabrielli:2024kbz , $\mu^{+}\mu^{-}$ Gao:2024leu , and $q\bar{q}$ Cheng:2025cuv . Some of these other states go beyond two qubits and probe qutrits, multipartite systems Sakurai:2023nsc ; Aguilar-Saavedra:2024whi ; Morales:2024jhj ; Subba:2024mnl , and other quantum informational concepts Altomonte:2023mug ; Barr:2024djo ; Altomonte:2024upf . Studies of the $\tau^{+}\tau^{-}$ system have been suggested at $e^{+}e^{-}$ colliders, including at the FCC-ee Altakach:2022ywa ; Fabbrichesi:2024wcd , at the CEPC Ma:2023yvd , at Belle-II Ehataht:2023zzt , and at the BEPC Han:2025ewp . [^1] See also Ref. larry .
Generally, $t\bar{t}$ has been considered the canonical qubit system in hadron colliders, whereas $\tau^{+}\tau^{-}$ has been considered the canonical qubit system in lepton colliders. Although both systems have at least two neutrinos in their final states, in the $t\bar{t}$ system there are enough kinematic constraints, save for discrete ambiguities, to solve the neutrino momenta at hadron or lepton colliders, while for $\tau^{+}\tau^{-}$ there are only enough constraints at lepton colliders due to the well-determined collision energies. Because of this difficulty, there has not been a study of $\tau^{+}\tau^{-}$ state at the LHC. [^2] Ref. Fabbrichesi:2022ovb presented theoretical-level results for $\tau^{+}\tau^{-}$ at the LHC but did not provide a practical way to measure the quantum state. See also Ref. LoChiatto:2024dmx . In this work, we leverage the power of ML to open up this notoriously difficult channel and show that it has excellent sensitivity to detecting quantum entanglement and Bell nonlocality, already with the LHC’s current dataset.
@@ 2.2 The Decay Method
The spin analyzing power varies from 1-1 to +1+1 corresponding to maximally anti-correlated and maximally correlated, respectively. A spin analyzing power of 0 indicates no correlation between the spin and the spin analyzer direction.
@@ 2.4 The $\tau^{+}\tau^{-}$ State
The quantum state of $\tau^{+}\tau^{-}$ in production $pp\to\tau^{+}\tau^{-}X$ has been discussed in Ref. Fabbrichesi:2022ovb (see also Banerjee:2023qjc ). Below the $Z$ boson mass $m_{Z}$ where the $Z$ contribution is negligible, the $\tau^{+}\tau^{-}$ state is dominantely produced through an $s$-channel photon leading to a triplet state that is nearly maximally entangled. As the $Z$ contribution grows and the interference with the $\gamma$ contribution becomes sizable, the mixing between the $Z$ and $\gamma$ leads to a nearly separable state as seen in Fig. 1. The minimum entanglement occurs near $m_{\tau\tau}\approx 70~{\rm GeV}$.
@@ 4 Neutrino Momentum Reconstruction Using a Generative Model
Compared with existing approaches, our method presents several notable advancements. Firstly, we introduce an additional Transformer layer within the generation head, thereby significantly enhancing the representational capability of the model. Secondly, we optimize the sampling procedure by replacing the modified Denoising Diffusion Probabilistic Model (DDPM) used in OmniLearn with the simpler Denoising Diffusion Implicit Model (DDIM), enabling increased sampling steps without incurring substantial computational costs. Thirdly, we incorporate random rotations into the training data augmentation process, greatly enlarging the dataset and enabling larger batch sizes, thus improving the training efficiency of the Transformer model while explicitly leveraging the rotational symmetry inherent in physical processes. Lastly, the nearest-neighbor calculations are dynamically tailored according to the number of visible components resulting from $\tau$ decays; specifically, we select the six nearest neighbors for decay channels involving $\rho$ particles and only two nearest neighbors for other channels, effectively capturing the distinct physical characteristics associated with each decay channel.
@@ 4.1 Model Architecture and Training Strategy
A key limitation of generative approaches in collider physics is their reliance on large Monte Carlo datasets. To mitigate this, we employ an efficient data augmentation strategy, where each event is randomly rotated in the transverse plane relative to the center-of-mass frame. This technique enhances generalization without requiring additional MC event generation, significantly reducing computational costs. By applying this augmentation, we increase the training statistics by approximately 50 times, reaching a range of 11M to 1010M effective events across all subchannels.
@@ 5.1 Event Selection and Categorization
[…]
$80~{\rm GeV}<m_{\tau\tau}<100~{\rm GeV},\qquad\qquad 0.6<2\theta_{\tau}/\pi<1.0.$ (19)
@@
Subchannel Prongness $|\kappa_{A}\times\kappa_{B}|$ $\tau\tau\to\ell\pi$ $\tau\tau\to\ell\rho$ $\tau\tau\to\pi\pi$ $\tau\tau\to\pi\rho$ $\tau\tau\to\rho\rho$
@@
$e\pi$ e - 1p0n 0.33 $1245.85\pm 3.25$ $80.33\pm 1.23$ $0.09\pm 0.02$ $14\pm 0.03 0.0$$01\pm 0.01$
$\mu\pi$ $\mu$ - 1p0n 0.34 $1640.89\pm 3.73$ $106.80\pm 1.42$ $04\pm 0.01 0.0$$5\pm 0.02 <0.0$1<0.01
$e\rho$ e - 1p1n 0.14 $25.32\pm 0.46$ $3603.19\pm 8.23$ <0.01<0.01 $0.13\pm 0.03$ $0.47\pm 0.08$
$\mu\rho$ $\mu$ - 1p1n 0.15 $36.45\pm 0.56$ $4869.69\pm 9.57$ <0.01<0.01 $0.09\pm 0.02$ $0.31\pm 0.06$
Signal Region ($80~{\rm GeV}<m_{\tau\tau}<100~{\rm GeV}$ & $0.6<2\theta_{\tau}/\pi<1)$
@@ 5.2 Backgrounds and Systematic Uncertainties
[…]
@@
$\pi\pi$ <0.01<0.01 $41\pm 0.29 <0.0$1<0.01 $96\pm 0.22 <0.0$1<0.01 $1.37\pm 0.37$
$\pi\rho$ $0.77\pm 0.55$ $1.45\pm 0.55$ <0.01<0.01 $3.28\pm 0.41$ <0.01<0.01 $5.50\pm 0.87$
$\rho\rho$ <0.01<0.01 $4.14\pm 0.93$ <0.01<0.01 $10.13\pm 0.71$ <0.01<0.01 $14.27\pm 1.17$
$e\pi$ $4.25\pm 1.28$ $3.31\pm 0.83$ $1.51\pm 0.75$ $42.33\pm 1.46$ <0.01<0.01 $51.40\pm 2.24$
$e\rho$ $28.20\pm 3.30$ $12.43\pm 1.60$ $2.64\pm 1.00$ $120.53\pm 2.46$ <0.01<0.01 $163.80\pm 4.53$
$\mu\pi$ $2.70\pm 1.02$ $5.39\pm 1.06$ <0.01<0.01 $50.89\pm 1.60$ <0.01<0.01 $58.98\pm 2.17$
$\mu\rho$ $34.00\pm 3.62$ $12.84\pm 1.63$ <0.01<0.01 <0.01<0.01 <0.01<0.01 $46.84\pm 3.97$
Signal Region ($80~{\rm GeV}<m_{\tau\tau}<100~{\rm GeV}$ & $0.6<2\theta_{\tau}/\pi<1)$
$\pi\pi$ <0.01<0.01 <0.01<0.01 <0.01<0.01 $05\pm 0.05 <0.0$1<0.01 $0.05\pm 0.05$
$\pi\rho$ <0.01<0.01 <0.01<0.01 <0.01<0.01 $20\pm 0.10 <0.0$1<0.01 $0.20\pm 0.10$
$\rho\rho$ <0.01<0.01 <0.01<0.01 <0.01<0.01 $96\pm 0.22 <0.0$1<0.01 $0.96\pm 0.22$
$e\pi$ <0.01<0.01 $21\pm 0.21 <0.0$1<0.01 $3.93\pm 0.45$ <0.01<0.01 $4.14\pm 0.49$
$e\rho$ $2.70\pm 1.02$ $62\pm 0.36 <0.0$1<0.01 $12.60\pm 0.80$ <0.01<0.01 $15.92\pm 1.34$
$\mu\pi$ <0.01<0.01 $83\pm 0.41 <0.0$1<0.01 $5.59\pm 0.53$ <0.01<0.01 $6.42\pm 0.67$
$\mu\rho$ $2.32\pm 0.95$ $1.04\pm 0.46$ <0.01<0.01 <0.01<0.01 <0.01<0.01 $3.35\pm 1.05$
@@ 5.2.1 Systematics Uncertainties
[…]
@@
Figure 4: Comparison of reconstructed neutrino kinematic distributions under systematic variations from soft MET along the $x$ direction for $\ell\rho$ channel. The three columns correspond to transverse momentum ($p_{T}$), pseudorapidity ($\eta$), and azimuthal angle ($\phi$). Each plot includes the nominal reconstruction (teal) along with systematically shifted variations: $1\sigma$ (goldenrod) and $1\sigma$ (rose) shifts applied to neutrinos originating from $\tau^{+}$. The middle panels show the ratio of systematically varied reconstructions and nominal values to the truth values, while the bottom panels contain two-dimensional correlation contour plots illustrating the linear dependency between the truth and systematically shifted predictions. The contour lines represent density levels, with numerical labels indicating the percentage of total data enclosed within each contour.
@@ 5.3 Trigger Considerations
The di-tau trigger [^3] HLT_tau35_medium1_tracktwo_tau25_medium1_tracktwo_L1TAU20IM_2TAU12IM is designed for events with two hadronically decaying $\tau$-leptons, covering the $\pi\pi$, $\pi\rho$, and $\rho\rho$ subchannels. This trigger requires the leading $\tau$ to have $p_{T}>35$ GeV and the subleading $\tau$ to have $p_{T}>25$ GeV. The tau + muon trigger [^4] HLT_mu14_tau25_medium1_tracktwo is applied to events in the $\mu\pi$ and $\mu\rho$ channels, requiring a $\tau$ with $p_{T}>25$ GeV and a muon with $p_{T}>14$ GeV. Additionally, the single-muon trigger [^5] HLT_mu26_ivarmedium is used, requiring a muon with $p_{T}>26$ GeV. For events involving electrons, the tau + electron trigger [^6] HLT_e17_lhmedium_nod0_ivarloose_tau25_medium1_tracktwo is applied, requiring a $\tau$ with $p_{T}>25$ GeV and an electron with $p_{T}>17$ GeV. The single electron trigger [^7] HLT_e26_lhtight_nod0_ivarloose is also utilized, selecting events with an electron satisfying $p_{T}>26$ GeV.
[…]